Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 211     Entry time: 28 May 2023, 22:57
Author: Martin Teichmann 
Group: SCS 
Subject: Shift summary all week 

we were trying to make THz and cross correlate this to the pump laser.

We managed to see THz and an auto-correlation with the pump laser therein. Unfortunately we did not see any THz generation using the FEL.

Problems:

Many. Only one problem was related to the FEL: weirdly, the zeroth order of the monochromator beam was much weaker than the first order, which is worrysome.

ELOG V3.1.4-7c3fd00